# GAIA: Rethinking Action Quality Assessment for AI-Generated Videos

### v1

@@ Abstract
Assessing action quality is both imperative and challenging due to its significant impact on the quality of AI-generated videos, further complicated by the inherently ambiguous nature of actions within AI-generated video (AIGV). Current action quality assessment (AQA) algorithms predominantly focus on actions from real specific scenarios and are pre-trained with normative action features, thus rendering them inapplicable in AIGVs. To address these problems, we construct GAIA, a Generic AI-generated Action dataset, by conducting a large-scale subjective evaluation from a novel causal reasoning-based perspective, resulting in 971,244 ratings among 9,180 video-action pairs. Based on GAIA, we evaluate a suite of popular text-to-video (T2V) models on their ability to generate visually rational actions, revealing their pros and cons on different categories of actions. We also extend GAIA as a testbed to benchmark the AQA capacity of existing automatic evaluation methods. Results show that traditional AQA methods, action-related metrics in recent T2V benchmarks, and mainstream video quality methods correlate poorly with human opinions, indicating a sizable gap between current models and human action perception patterns in AIGVs. Our findings underscore the significance of action quality as a unique perspective for studying AIGVs and can catalyze progress towards methods with enhanced capacities for AQA in AIGVs.
@@ 1 Introduction
However, there is a significant gap in the existing AQA research. First, prior work has contributed multiple AQA datasets, which predominantly focus on domain-specific actions from real videos and collect coarse-grained expert-only human ratings [71, 112, 70] on limited dimensions. Meanwhile, the content discrepancies in those AQA videos are really subtle, since the action subjects are usually performing similar actions in a fixed environment, such as swimming and diving in the natatorium, or gymnastics in the gym, lacking the consideration of scene diversity. Second, the existing AQA approaches mainly follow a pose-based or vision-based feature extraction, aggregation, and score regression ternary form, which usually adopt powerful 3D backbone networks that are pre-trained on large action recognition datasets [13, 87] for better feature migration. Nevertheless, a distinguishing characteristic of generated videos is that they may contain atypical actions with various body or object artifacts over time [21, 63], such as aberrant limb count, irrational object shape, and physically implausible motion, due to the stochasticity and unstable nature of the diffusion process. In such cases, the model learned from real action videos may fail in AIGVs with worse prediction performance. At present, it remains unclear to what degree any T2V model can achieve visually rational action generation that varies in action categories, much less the cognitive mechanism of action quality that affects human perception.
@@
To address these issues, we present GAIA, a Generic AI-generated Action dataset encompassing 9,180 AI-generated videos from 18 T2V models, spanning both lab studies and commercial platforms, which covers a variety of whole-body, hand, and facial actions. Specifically, we recruit 54 participants and conduct a large-scale human evaluation to evaluate the action quality first-of-this-kind from three causal reasoning-based perspectives: subject quality, action completeness, and action-scene interaction. Crucially, it provides quantifiable action state estimations based on the behavior of human reasoning in perceiving an action. In theory, this makes complicated coupled action quality approachable and tractable. In practice, the full potential of multi-dimension methods remains largely untapped due to a scarcity of existing datasets, exacerbated by the difficulty of obtaining reliable group subjective opinions. This complements earlier research, which primarily concentrated on AQA under a single real scenario and lacked rating granularity. We demonstrate the value and type of insights GAIA enables by using it to evaluate the action generation ability and weaknesses across different categories of 18 representative T2V models (several times more than existing benchmarks [21, 46, 63, 61]). Moreover, we contribute a holistic benchmark based on GAIA, which reveals that the existing AQA methods and action-related automatic metrics even video quality assessment (VQA) approaches, correlate poorly with human evaluation. Overall, our study could serve as a pilot for future endeavors aimed at developing accurate AQA methods in generative scenarios while providing substantial insights for better defining the quality of AIGVs.
@@ 2 Related Work
Action Quality Assessment. Action quality assessment, which aims to discriminate and evaluate how well an action is performed, has been widely explored in applications such as sports events [88, 74, 73, 71, 72, 66, 52], healthcare [114, 32, 101, 70, 24], and public security [35, 89, 93]. Early works [74, 98] solely consider human pose-based features while neglecting the relations among joints and other action quality-related visual cues (e.g. splash in diving or barbell position in weightlifting). Later, researchers introduced graph structure to model the joint motion information spatially and temporally [69, 11, 27, 115]. In another line, vision-based approaches [73, 107, 108, 115, 117, 24] combined 3D convolutional neural networks (C3D) [94] with context-aware modules to extract motion-oriented spatial-temporal visual features for assessing action quality. As shown in Tab. 1, there are numerous datasets for AQA, which predominantly focus on the sports domain from real scenarios, while the problem of considering action quality in AI-generated videos remains unexplored.
@@
Evaluations on Video Generative Models. Early video generation models shared the same frame-wise evaluation metrics as T2I models, such as Inception Score (IS) [80], Fréchet Inception Distance (FID) [40], and CLIPScore [75], as well as their variants for video [96, 97, 79]. These metrics are all group-targeted and not suitable for assessing a single video. For text-to-video (T2V) models, several benchmarks [21, 61, 46, 63, 51] have been proposed to assess various perspectives like video fidelity [51], temporal quality [61], text-video alignment [46, 63]. Despite covering various dimensions, these works lack specificity and breadth with limited model exploration and human group annotation.
@@ 3.2 Task Definition: the Action Syllogism
Considering the peculiar characteristics of AI-generated videos, to collect a more explainable and nuanced understanding of public perception on action assessment, instead of collecting professional skill scores as in existing AQA studies [70, 108], we opt to collect annotations from a novel perspective, namely the causal reasoning syllogism [86]. Specifically, we decompose an action process into three parts: 1) action subject as major premise, 2) action completeness as minor premise, and 3) interaction between action and scenes as conclusion, according to the syllogism theory. The rationales for this strategy are as follows: (a) As the visual saliency information in action-oriented videos, the rendering quality of the action subject can profoundly affect the visibility of the action, while humans excel at perceiving such generated artifacts [19, 46, 64]. (b) Moreover, unlike parallel-form feedbacks, the order of these three parts in action syllogism inherently aligns with the human reasoning process. For instance, while human annotators are shown with an action scene about “A musician is playing the piano”, they can intuitively reasoning like: ❶ a musician as the major premise, which is the subject to execute the action of playing the piano; ❷ appearance or completeness of action as the minor premise, containing the spatial and temporal boundaries in the given scenario; ❸ phenomenon of the keys being pressed as the conclusion describing a reasonable result considering both constraints. This reasoning-form evaluation can not only facilitate the community to improve the shortcomings of generative models to eventually produce plausible videos but also can be transformed into learning objectives to design more subjectively consistent objective quality assessment algorithms.
@@ 3.3 Subjective Action Quality Assessment
Main Process. We adopted a single-stimulus methodology in this evaluation and asked participants to focus on the given action keyword as well as the corresponding prompt and evaluate three action-related dimensions of AI-generated videos, i.e., subject quality, action completeness, and action-scene interaction, by dragging the slide button at a $[0,100]$ continuous rating scale. We randomly divided the 9,180 videos in GAIA into 31 sessions, with each session, except the last, comprising 300 videos. Ten golden videos with expert opinions from the real-world action database [14] were added to each session as an inspection to control the scoring deviations. Only participants who had a high agreement (Pearson linear correlation coefficient, $PLCC>0.7$) with the mean opinion score (MOS) from experts were eligible to continue to the next session, leaving 48 remaining. To reduce visual fatigue, there is a rest segment with at least 15 minutes per 150 videos [83, 19, 18].
@@
In summary, it took participants approximately 2.6 hours to finish one session, and all experiments took over a month to complete. Each participant was compensated \$12 for each session according to the current ethical standard [85].
@@ 3.4 Dataset Statistics and Analysis
Overall Observations. Each data sample in GAIA consists of four elements: the action keyword $k$, the corresponding prompt $t$, the generated video $v$, and the action quality-related human annotations {${{\text{\{MO}}{{\text{S}}_{a}}\}_{a\in\mathcal{A}}}$. $\mathcal{A}$ is the collection of three perspectives. Fig. 4 illustrates two examples of generated videos with small (shaking hands) and large (riding bike) movements. In Fig. 1(c), we visualize the 3D scatter map of human-annotated subject quality, action completeness, and action-scene interaction scores in the GAIA dataset and examine three extreme cases, where two dimensions most differently or consistently (noted in purple circles). In general, the generated videos receive lower-than-average human ratings ($\mu_{s}=35.48$, $\mu_{c}=33.81$, $\mu_{i}=30.25$) on three perspectives, suggesting the inferior performance of existing models to produce artifact-free videos with coherent actions.
@@
From Tab. 4, we notice a significantly higher correlation between $\mathrm{MOS}_{c}$ and $\mathrm{MOS}_{i}$ (0.931 Spearman’s $\rho$, 0.791 Kendall’s $\tau$) than other pairs, indicating that action completeness is a great premise of its rich interaction with the scene context, which further demonstrates the syllogism-based action evaluation strategy.
Model-wise Comparison. As illustrated in Fig. 3 and Fig. 5, the commercial T2V models generally perform better than models from lab studies in three evaluated dimensions. Most models exhibit left-skewed MOS distribution in all three dimensions. Among them, VideoCrafter2 [16] and Morph Studio [4] are basically the best models in their respective fields (see Fig. 13 in the Appendix for detailed ranking in all dimensions). Additionally, we can observe a trend of increasing performance year by year, from the Text2Video-zero [48] and ModelScope [99] released in March 2023 to the VideoCrafter2 [16] in early 2024. Nevertheless, most models prove decent proficiency on one single dimension, i.e., better subject quality than action completeness and action-scene interaction, which exposes the defects of existing models in producing temporal coherent and complete actions.
@@
Surprisingly, the newly proposed Mora [111] significantly underperforms other models in all three perspectives, we speculate that it is limited by the core dependency model in its demo code, stable video diffusion (SVD), an earlier image-to-video model. Furthermore, comparing the two resolution versions of VideoCrafter1 ($512\times 320$ and $1024\times 576$) [15], as well as the commercial models and the lab studies (with an average resolution of $596\times 378$ and $1385\times 835$), we can conclude that higher resolution plays an important role in improving action recognizability, resulting in advancements in the subject quality and action completeness. A similar conclusion applies to the performance gains as the frame rate increases.
Class-wise Comparison. We investigate the MOS distribution across action categories via box plots, as presented in Fig. 6. It can be observed that the $\mathrm{MOS_{s}}$, $\mathrm{MOS_{c}}$, and $\mathrm{MOS_{i}}$ of complex actions such as jumping/throwing and racquet-bat are lower than actions with small movements (e.g., communication, touching person, and using tools), indicating that existing T2V models struggle to render actions with drastic motion changes, where atypical body postures are more easily involved. Additionally, when it comes to the local hand action categories, the actions contain subtle movements, e.g., rotate or move fingers/palm, or numeral representation receive significantly lower MOSs than others, showing the inferior capacity of generating fine-grained actions. Specifically, the frequency of outliers in Fig. 6 reflects the response variance of evaluated models under specific action word conditions, which further supports the above viewpoints. Beyond the above observations, we further analyze the diversity of contents in GAIA (see Sec. B.2).
@@ 4.2 Main Results and Analysis
Do conventional AQA methods still work? As shown in Tab. 5, all AQA methods perform poorly with an average SRCC of $0.4367$, $0.4722$, and $0.4664$ in terms of subject quality, action completeness, and action-scene interaction, respectively. Specifically, USDL takes a manually defined Gaussian distribution as the learning objective to address the uncertainty during the human assessment process, which, on the contrary, exacerbates the prediction inaccuracy. Benefiting from the dynamic-static hybrid stream, ACTION-NET can capture the body postures at specific moments during an action process and thus performs marginally better than the rest models in terms of subject quality. For CoRe and TSA, the input requirement is a pairwise query and exemplary video, which is not exactly applicable to AIGVs, since the same action from different models can vary significantly from generation quality to content scenarios, failing the contrastive regression strategy [112, 108]. Most importantly, plagued by the generation quality of AIGV itself, it is difficult for those commonly used inflated 3D ConvNets (I3D) backbone to learn normative action features as in Kinetics [14]. In general, existing AQA methods focus mainly on assessing actions in a similar environment, where the differences between videos are subtle, which is in accord with its goal (most for specific tasks rather than a generic AQA). This also highlights the necessity of our proposed GAIA dataset.
@@
Comparison to the VQA methods. Considering the intrinsic correlation of action quality on the content quality of videos, we select seven representative VQA methods to validate whether VQA approaches are applicable for AQA tasks in AI-generated scenarios, as shown in the third part of Tab. 5. We can observe that VQA methods surpass all AQA methods and action-related metrics by a large margin (on average $25.04\%$ and $131.1\%$ better than their respective best methods in terms of SRCC) in the subject quality dimension, while deep learning-based VQA methods perform better than traditional VQA methods (TLVQM and VIDEVAL) that rely on handcraft features. Notably, all VQA methods exhibit a relatively superior capacity to evaluate the subject quality than assessing the action completeness and action-scene interaction, indicating a potential emphasis on low-level technical distortions such as noises, sharpness, blur, and artifacts within the current VQA frameworks, which may not fully encapsulate the temporal-level normativity and interactive facets of action content. Such a conclusion is also supported by evidence from being equipped with different quality-aware initializations, as BVQA and DOVER are pre-trained with spatial distortion-dominated datasets [22, 33, 45, 30, 109]. Moreover, BVQA and SimpleVQA utilize the SlowFast [31] model as the motion feature extractor, which has proven effective in action recognition tasks benefit by its slow and fast pathway designs in learning both spatial semantic and motion information, nonetheless, faces constraints due to the limited frames of AIGVs. Another plausible explanation for these subpar performances is the pure regression-based prediction strategy that lacks consideration of textual information, as the same MOS for different actions could lead to a large visual discrepancy.
@@
Evaluation on video-text alignment metrics. We further evaluate the performance of video-text alignment metrics in measuring action quality considering their capacity in cross-modality feature mapping. Specifically, we compute the cosine similarity between the image embedding and the action prompt embedding to record a deviation degree between the sketch of the content and target action semantics. As listed in Tab. 5, the widely used CLIPScore achieves a weak correlation with human opinion, especially in the subject quality dimension, while performing relatively better with respect to action completeness and action-scene interaction dimensions. We conjecture that this is because such alignment-based metrics are intrinsically sensitive to high-level vision information (action semantics) rather than low-level generative flaws (e.g., blur, noise, textures). Meanwhile, we see a decent performance gain on evaluated dimensions ($+8.2\%$, $+9.6\%$, $+10.9\%$, $+13.1\%$ in terms of SRCC) when replacing CLIP with a more powerful VLM, such as InternLM-XComposer2-VL, showing an underlying possibility of building more accurate AQA metrics as VLMs evolve.

### v2

@@ Abstract
Assessing action quality is both imperative and challenging due to its significant impact on the quality of AI-generated videos, further complicated by the inherently ambiguous nature of actions within AI-generated video (AIGV). Current action quality assessment (AQA) algorithms predominantly focus on actions from real specific scenarios and are pre-trained with normative action features, thus rendering them inapplicable in AIGVs. To address these problems, we construct GAIA, a Generic AI-generated Action dataset, by conducting a large-scale subjective evaluation from a novel causal reasoning-based perspective, resulting in 971,244 ratings among 9,180 video-action pairs. Based on GAIA, we evaluate a suite of popular text-to-video (T2V) models on their ability to generate visually rational actions, revealing their pros and cons on different categories of actions. We also extend GAIA as a testbed to benchmark the AQA capacity of existing automatic evaluation methods. Results show that traditional AQA methods, action-related metrics in recent T2V benchmarks, and mainstream video quality methods perform poorly with an average SRCC of 0.454, 0.191, and 0.519, respectively, indicating a sizable gap between current models and human action perception patterns in AIGVs. Our findings underscore the significance of action quality as a unique perspective for studying AIGVs and can catalyze progress towards methods with enhanced capacities for AQA in AIGVs.
@@ 1 Introduction
However, there is a significant gap in the existing AQA research. First, prior work has contributed multiple AQA datasets, which predominantly focus on domain-specific actions from real videos and collect coarse-grained expert-only human ratings [76, 118, 75] on limited dimensions. Meanwhile, the content discrepancies in those AQA videos are often subtle, as the action subjects typically perform similar actions within a consistent environment. Examples include swimming and diving in a natatorium or gymnastics in a gym, which lacks consideration for scene diversity. Second, the existing AQA approaches mainly follow a pose-based or vision-based feature extraction, aggregation, and score regression ternary form, which usually adopt powerful 3D backbone networks that are pre-trained on large action recognition datasets [13, 93] for better feature migration. Nevertheless, a distinguishing characteristic of generated videos is that they may contain atypical actions with various body or object artifacts over time [22, 68], such as aberrant limb count, irrational object shape, and physically implausible motion, due to the stochasticity and unstable nature of the diffusion process. In such cases, the model learned from real action videos may fail in AIGVs with worse prediction performance. At present, it remains unclear to what degree any T2V model can achieve visually rational action generation that varies in action categories, much less the cognitive mechanism of action quality that affects human perception.
@@
To address these issues, we present GAIA, a Generic AI-generated Action dataset encompassing 9,180 AI-generated videos from 18 T2V models, spanning both lab studies and commercial platforms, which covers a variety of whole-body, hand, and facial actions. Specifically, we recruit 54 participants and conduct a large-scale human evaluation to evaluate the action quality first-of-this-kind from three causal reasoning-based perspectives: subject quality, action completeness, and action-scene interaction. Among them, as the major premise of an action, the quality of the action subject directly affects the whole action process. Assessing action completeness ensures that the generated action is not only temporally coherent but also logically and narratively complete. Action-scene interaction considers the spatial relationships, environmental factors, and interactions with other elements within the scene that can influence the perception of the action’s quality and realism. Crucially, it provides quantifiable action state estimations based on the behavior of human reasoning in perceiving an action. In theory, this makes complicated coupled action quality approachable and tractable. In practice, the full potential of multi-dimension methods remains largely untapped due to a scarcity of existing datasets, exacerbated by the difficulty of obtaining reliable group subjective opinions. This complements earlier research, which primarily concentrated on AQA under a single real scenario and lacked rating granularity.
We prove the value and type of insights GAIA enables by using it to evaluate the action generation ability and weaknesses across different categories of 18 representative T2V models (several times more than existing benchmarks [22, 50, 68, 66]). Moreover, we contribute a holistic benchmark based on GAIA, which reveals that the existing AQA methods and action-related automatic metrics even video quality assessment (VQA) approaches, correlate poorly with human evaluation. Overall, our study could serve as a pilot for future endeavors aimed at developing accurate AQA methods in generative scenarios while providing substantial insights for better defining the quality of AIGVs.
@@ 2 Related Work
Action Quality Assessment. Action quality assessment, which aims to discriminate and evaluate how well an action is performed, has been widely explored in applications such as sports events [94, 79, 78, 76, 77, 71, 57], healthcare [120, 34, 107, 75, 26], and public security [39, 95, 99]. Early works [79, 104] solely consider human pose-based features while neglecting the relations among joints and other action quality-related visual cues (e.g. splash in diving or barbell position in weightlifting). Later, researchers introduced graph structure to model the joint motion information spatially and temporally [74, 11, 29, 121]. In another line, vision-based approaches [78, 113, 114, 121, 123, 26] combined 3D convolutional neural networks (C3D) [100] with context-aware modules to extract motion-oriented spatial-temporal visual features for assessing action quality. As shown in Tab. 1, there are numerous datasets for AQA, which predominantly focus on the sports domain from real scenarios, while the problem of considering action quality in AI-generated videos remains unexplored. Furthermore, human activities often occur in specific scene contexts, e.g., swimming in a swimming pool. However, it is also possible to enjoy champagne in the pool. Training an action quality assessment model for more diverse AI-generated actions using existing video datasets thus inevitably introduces such bias, which may opposed to paying attention to the actual action in the scene. Choi et al. [23] proposed to mitigate scene bias by adding an adversarial loss for scene types and masking out the human actors. Similar operations include extracting the foreground and background parts of the video as data augmented pairs to improve the accuracy of action recognition [37]. Apart from pixel space augmentation, Gorpincenko et al. [36] extended it further to utilize the time domain to perform deeper levels of temporal perturbations, thus improving the robustness of action classifiers. The above studies illustrate the necessity of disentangling action process from scenes or decomposing the action process to mitigate the effect of representation bias.
From the perspective of data provenance, virtual worlds and game engines are plausible techniques to generate editable actions as synthetic data before the Generative AI Era. Such synthetic data has been used to train visual models for lots of computer vision tasks (e.g., object detection, recognition, pose estimation, and scene understanding) to extract visual priors. Desouza et al. [81] proposed a diverse, realistic, and physically plausible dataset of human action videos using virtual world simulation software. Experiments show that mixing both synthetic and real samples at the mini-batch level during training can significantly improve action recognition accuracy. Similarly, the controllability of both the type and quality of AI-generated actions (using different types of prompts and video generation models) offers a feasible solution for constructing large-scale action quality assessment datasets, especially for those irrational scenarios.
@@
Evaluations on Video Generative Models. Early video generation models shared the same frame-wise evaluation metrics as T2I models, such as Inception Score (IS) [86], Fréchet Inception Distance (FID) [44], and CLIPScore [80], as well as their variants for video [102, 103, 85]. These metrics are all group-targeted and not suitable for assessing a single video. For text-to-video (T2V) models, several benchmarks [22, 66, 50, 68, 56] have been proposed to assess various perspectives like video fidelity [56], temporal quality [66], text-video alignment [50, 68]. Despite covering various dimensions, these works lack specificity and breadth with limited model exploration and human group annotation. Our work differs from current research in three key aspects: 1) We created 510 distinct action prompts covering both coarse-grained and fine-grained actions, each applied with 18 T2V models for extensive assessment. 2) Our casual reasoning-based and multi-dimensional action quality evaluation offers valuable and comprehensive insights into video generation. 3) We have quantitatively validated a large amount of existing metrics that none of them performs well on the AI-generated action quality assessment task.
@@ 3.2 Task Definition: the Action Syllogism
Considering the peculiar characteristics of AI-generated videos, to collect a more explainable and nuanced understanding of public perception on action assessment, instead of collecting professional skill scores as in existing AQA studies [75, 114], we opt to collect annotations from a novel perspective, namely the causal reasoning syllogism [92, 53]. Specifically, we decompose an action process into three parts: 1) action subject as major premise, 2) action completeness as minor premise, and 3) interaction between action and scenes as conclusion, according to the syllogism theory. The rationale for this strategy is as follows: (a) The visibility of the action in videos is greatly affected by the rendering quality of the action subject, which is a crucial element of visual saliency information, while humans excel at perceiving such generated artifacts [20, 50, 69]. (b) Moreover, unlike parallel-form feedbacks, the order of these three parts in action syllogism inherently aligns with the human reasoning process. For instance, while human annotators are shown with an action scene about “A musician is playing the piano”, they can intuitively reason like: ❶ a musician as the major premise, which is the subject to execute the action of playing the piano; ❷ appearance or completeness of action as the minor premise, containing the spatial and temporal boundaries in the given scenario; ❸ phenomenon of the keys being pressed as the conclusion describing a reasonable result considering both constraints. This reasoning-form evaluation has many merits. First, by breaking down an action into its constituent parts, researchers can more clearly identify and analyze the specific elements that contribute to the perceived quality of the action. Second, such causal reasoning-based strategy is inherently aligned with human perception and can help in understanding how different parts of action are perceived by the public, which can lead to insights into what makes AI-generated action convincing or unconvincing. Third, this scheme allows for a comparative analysis of AI-generated action against natural human action, revealing where AI excels and where it may need improvement.
@@ 3.3 Subjective Action Quality Assessment
Main Process. We adopted a single-stimulus methodology in this evaluation and asked participants to focus on the given action keyword as well as the corresponding prompt and evaluate three action-related dimensions of AI-generated videos, i.e., subject quality, action completeness, and action-scene interaction, by dragging the slide button at a $[0,100]$ continuous rating scale. We randomly divided the 9,180 videos in GAIA into 31 sessions, with each session, except the last, comprising 300 videos. Ten golden videos with expert opinions from the real-world action database [14] were added to each session as an inspection to control the scoring deviations. Only participants who had a high agreement (Pearson linear correlation coefficient, $PLCC>0.7$) with the mean opinion score (MOS) from experts were eligible to continue to the next session, leaving 48 remaining. To reduce visual fatigue, there is a rest segment with at least 15 minutes per 150 videos [89, 20, 19]. In summary, it took participants approximately 2.6 hours to finish one session, and all experiments took over a month to complete. Each participant was compensated \$12 for each session according to the current ethical standard [91].
@@ 3.4 Dataset Statistics and Analysis
Overall Observations. Each data sample in GAIA consists of four elements: the action keyword $k$, the corresponding prompt $t$, the generated video $v$, and the action quality-related human annotations {${{\text{\{MO}}{{\text{S}}_{a}}\}_{a\in\mathcal{A}}}$. $\mathcal{A}$ is the collection of three perspectives. Fig. 4 illustrates two examples of generated videos with small (shaking hands) and large (riding bike) movements. In Fig. 1(c), we visualize the 3D scatter map of human-annotated subject quality, action completeness, and action-scene interaction scores in the GAIA dataset and examine three extreme cases, where two dimensions are most differently or consistently (noted in purple circles). In general, the generated videos receive lower-than-average human ratings ($\mu_{s}=35.48$, $\mu_{c}=33.81$, $\mu_{i}=30.25$) on three perspectives, suggesting the inferior performance of existing models to produce artifact-free videos with coherent actions. From Tab. 4, we notice a significantly higher correlation between $\mathrm{MOS}_{c}$ and $\mathrm{MOS}_{i}$ (0.931 Spearman’s $\rho$, 0.791 Kendall’s $\tau$) than other pairs, indicating that action completeness is a great premise of its rich interaction with the scene context, which further demonstrates the syllogism-based action evaluation strategy. More results are in Sec. B.2.1.
Model-wise Comparison. As illustrated in Fig. 3 and Fig. 6, the commercial T2V models generally perform better than models from lab studies in three evaluated dimensions. Most models exhibit left-skewed MOS distribution in all three dimensions. Among them, VideoCrafter2 [16] and Morph Studio [4] are basically the best models in their respective fields (see Fig. 13 in the Appendix for detailed ranking in all dimensions). Additionally, we can observe a trend of increasing performance year by year, from the Text2Video-zero [52] and ModelScope [105] released in March 2023 to the VideoCrafter2 [16] in early 2024. Nevertheless, most models prove decent proficiency on one single dimension, i.e., better subject quality than action completeness and action-scene interaction, which exposes the defects of existing models in producing temporal coherent and complete actions. Surprisingly, the newly proposed Mora [117] significantly underperforms other models in all three perspectives, we speculate that it is limited by the core dependency model in its demo code, stable video diffusion (SVD), an earlier image-to-video model. Furthermore, comparing the two resolution versions of VideoCrafter1 ($512\times 320$ and $1024\times 576$) [15], as well as the commercial models and the lab studies (with an average resolution of $596\times 378$ and $1385\times 835$), we can conclude that higher resolution plays an important role in improving action recognizability, resulting in advancements in the subject quality and action completeness. A similar conclusion applies to the performance gains as the frame rate increases.
Class-wise Comparison. We investigate the MOS distribution across action categories via box plots, as presented in Fig. 5. It can be observed that the $\mathrm{MOS_{s}}$, $\mathrm{MOS_{c}}$, and $\mathrm{MOS_{i}}$ of complex actions such as jumping/throwing and racquet-bat are lower than actions with small movements (e.g., communication, touching person, and using tools) ($p<0.01$, Two-side T-test), indicating that existing T2V models struggle to render actions with drastic motion changes, where atypical body postures are more easily involved. Additionally, when it comes to the local hand action categories, the actions contain subtle movements, e.g., rotate or move fingers/palm, or numeral representation receive significantly lower MOSs than others, showing the inferior capacity of generating fine-grained actions. Specifically, the frequency of outliers in Fig. 5 reflects the response variance of evaluated models under specific action word conditions, which further supports the above viewpoints. Beyond the above observations, we further analyze the diversity of contents in GAIA (see Sec. B.2).
@@ 4.2 Main Results and Analysis
Do conventional AQA methods still work? As shown in Tab. 5, all AQA methods perform poorly with an average SRCC of $0.4367$, $0.4722$, and $0.4664$ in terms of subject quality, action completeness, and action-scene interaction, respectively. Specifically, USDL takes a manually defined Gaussian distribution as the learning objective to address the uncertainty during the human assessment process, which, on the contrary, exacerbates the prediction inaccuracy. Benefiting from the dynamic-static hybrid stream, ACTION-NET can capture the body postures at specific moments during an action process and thus performs marginally better than the rest models in terms of subject quality. For CoRe and TSA, the input requirement is a pairwise query and exemplary video, which is not exactly applicable to AIGVs, since the same action from different models can vary significantly from generation quality to content scenarios, failing the contrastive regression strategy [118, 114]. Most importantly, plagued by the generation quality of AIGV itself, it is difficult for those commonly used inflated 3D ConvNets (I3D) backbone to learn normative action features as in Kinetics [14]. In general, existing AQA methods focus mainly on assessing actions in a similar environment, where the differences between videos are subtle, which is in accord with its goal (most for specific tasks rather than a generic AQA).
@@
Comparison to the VQA methods. Considering the intrinsic correlation of action quality on the content quality of videos, we select seven representative VQA methods to validate whether VQA approaches are applicable for AQA tasks in AI-generated scenarios, as shown in the third part of Tab. 5. We can observe that VQA methods surpass all AQA methods and action-related metrics by a large margin (on average $25.04\%$ and $131.1\%$ better than their respective best methods in terms of SRCC) in the subject quality dimension, while deep learning-based VQA methods perform better than traditional VQA methods (TLVQM and VIDEVAL) that rely on handcraft features. Notably, all VQA methods exhibit a relatively superior capacity to evaluate the subject quality than assessing the action completeness and action-scene interaction, indicating a potential emphasis on low-level technical distortions such as noises, sharpness, blur, and artifacts within the current VQA frameworks, which may not fully encapsulate the temporal-level normativity and interactive facets of action content. Such a conclusion is also supported by evidence from being equipped with different quality-aware initializations, as BVQA and DOVER are pre-trained with spatial distortion-dominated datasets [24, 35, 49, 32, 115]. Moreover, BVQA and SimpleVQA leverage the SlowFast model [33] as their motion feature extractor. This model has demonstrated effectiveness in various action recognition tasks due to its dual pathway design, which captures both spatial semantics and motion information parallelly. However, it encounters problems when applied to AIGVs, primarily because of the limited frames. Another plausible explanation for these subpar performances is the pure regression-based prediction strategy that lacks consideration of textual information, as the same MOS for different actions could lead to a large visual discrepancy.
@@
Evaluation on video-text alignment metrics. We further evaluate the performance of video-text alignment metrics in measuring action quality considering their capacity in cross-modality feature mapping. Specifically, we compute the cosine similarity between the image embedding and the action prompt embedding to record a deviation degree between the sketch of the content and target action semantics. As listed in Tab. 5, the widely used CLIPScore achieves a weak correlation with human opinion, especially in the subject quality dimension, while performing relatively better with respect to action completeness and action-scene interaction dimensions. We conjecture that this is because such alignment-based metrics are intrinsically sensitive to high-level vision information (action semantics) rather than low-level generative flaws (e.g., blur, noise, textures). Meanwhile, we see a decent performance gain on evaluated dimensions ($+8.2\%$, $+9.6\%$, $+10.9\%$, $+13.1\%$ in terms of SRCC) when replacing CLIP with a more powerful VLM, such as InternLM-XComposer2-VL, showing an underlying possibility of building more accurate AQA metrics as VLMs evolve. We also conduct a T-test with a 95% confidence level to assess the statistical significance of the performance difference between any two methods (Tab. 9 and Fig. 15). More results are discussed in Sec. D.